OpenStudy (anonymous):

Sequence 2,6,18,54 Complete each statement? A(1) = 2 =2x3 And then there is a little empty box next to the 3 to fill in! please help and explain what i do?

OpenStudy (hoblos):

2=2x3^0 A(1) = 2x3^0 A(2)=2x3^1 A(3)=2x3^2 A(4)=2x3^3 . .

OpenStudy (hoblos):

do you get it?